A 60 g solid (1 j\gk) at 30 degree celcius is submerge to 100 mL water having an initial temperature of 80 degree cecius. The equilibrium temperature is 73.729 degree celcius. Ignore the heat effects of the water container. What is the specific heat capacity of dihydrogen monoxide?
1
Expert's answer
2019-10-07T07:21:59-0400

First, we should write the heat balance equation, which is based on the energy conservation law (the heat amount taken from water will be equal to the heat amount taken by the solid):


"m_{H_2O}C_{H_2O}(t_{H_2O}-t_e)=m_sC_s(t_e-t_s)"


or,


"C_{H_2O}=\frac{m_sC_s(t_e-t_s)}{m_{H_2O}(t_{H_2O}-t_e)}"


Now, it is the right time to put all the data in:


"C_{H_2O}=\frac{60g*1Jg^{-1}K^{-1}(73.729C-30.000C)}{100g*(80.000C-73.729C)}=4.184Jg^{-1}K^{-1}"


This value found is the same as a widely known water heat capacity constant, which could be found in practically each handbook.